OpenStudy (anonymous):

the square root of 1+\[1/(\sqrt{2}+1) +1/(\sqrt{3} +\sqrt{2}) + 1/(\sqrt{4} +\sqrt{3}) +............... 1/(\sqrt{324} +\sqrt{323}) is\]

OpenStudy (anonymous):

options are 3root2 1/root2 2root3

OpenStudy (anonymous):

the answer is 3root2

OpenStudy (anonymous):

do you mean square root of this whole term written in long?

OpenStudy (anonymous):

yes we have to find the square root

OpenStudy (anonymous):

i think there is some trick.........

OpenStudy (anonymous):

yes it is i am giving u the hints.. rationalize the each terms..!

OpenStudy (anonymous):

the square root of 1+\[1/(\sqrt{2}+1) +1/(\sqrt{3} +\sqrt{2}) + 1/(\sqrt{4} +\sqrt{3}) +............... 1/(\sqrt{324} +\sqrt{323}) is\]

OpenStudy (anonymous):

this is the question plzzz look at it

OpenStudy (anonymous):

he square root of 1+ 1/(2√+1)+1/(3√+2√)+1/(4√+3√)+...............1/(324−−−√+323−−−√)is

OpenStudy (anonymous):

dont forgot to add that 1

OpenStudy (amistre64):

do you know what a conjugate of a number is?

OpenStudy (amistre64):

\[\frac{1}{a+b}*\frac{a-b}{a-b}=\frac{a-b}{a^2-b^2}\]

OpenStudy (anonymous):

rationalizing the terms in denomiantor of each terms u will get 1.. & in numerator the terms like this.. \[\sqrt{2}-1+\sqrt{3}-\sqrt{2}+.........................\sqrt{324}-\sqrt{323}\] so each terms get cancelled and we are left with \[\sqrt{324}\]

OpenStudy (amistre64):

that called a telescoping sum right ?

OpenStudy (anonymous):

take the square root of sqrt(324) that is square root of 324 which is 3root2

OpenStudy (anonymous):

oh thanzzzz u really is a worth for me 1000000 thanzzzz